OpenStudy (erinweeks):

H E L P ! The area of a rectangle is x2 - 2x - 15 and the length of the rectangle is x + 3. Find the width of the rectangle.

OpenStudy (erinweeks):

answers A. x – 5 B. x2 + 15 C. x + 5 D. (x – 2)(x + 3)

OpenStudy (valpey):

Area = Length * Width \[x^2-2x-15 = (x+3) * Width=(x+3)(x\pm ?)\]

OpenStudy (erinweeks):

do you know the answer or how to help more ?

OpenStudy (erinweeks):

????

OpenStudy (anonymous):

factor the left hand side of the equation

OpenStudy (erinweeks):

i dont know how to ..

OpenStudy (anonymous):

have you done it ever before? factoring trinomials?

OpenStudy (anonymous):

or you can do long division

OpenStudy (erinweeks):

not really , no !

OpenStudy (erinweeks):

i guessed i was D but it was wrong

OpenStudy (valpey):

\[x^2−2x−15=(x+3)∗Width=(x+3)(x+A)\]\[x^2−2x−15=(x*x)+ (x*A)+(3*x)+ (3*A)=x^2+Ax +3x+3A\]\[x^2−2x−15=x^2+Ax+3x+3A=x^2+(3+A)x+3A\]\[x^2−2x−15=x^2+(3+A)x+3A\]Solution where -2 = (3+A) and -15 = 3A